# Fixing It in Post: A Comparative Study of LLM Post-Training Data Quality and Model Performance

### v2

@@ B.1 Tulu
The Tulu dataset was created to bridge proprietary and open-source post-training data by leveraging publicly available datasets, persona-driven synthetic prompts, and rigorous decontamination procedures to mitigate test set leakage. Specifically, Lambert et al. [7] collected 23,327,961 candidate prompts from over 20 distinct sources, curating a multi-skill SFT corpus that comprises 939,344 samples, forming the original Tulu-3-SFT-Mix data mixture. The Tulu subsets and their respective samples can be broadly categorized into nine high-level groups:
  • General: OpenAssistant (OASST1) [19], No Robots [51], WildChat [8], UltraFeedback (Tülu HC-10) [47]
  • Knowledge Recall: FLAN v2 [52], SciRIFF [53], TableGPT [54]
  • Math: Persona MATH, Persona MATH (Grade) [55]
  • Reasoning: Persona Algebra [56], OpenMathInstruct2 [10], NuminaMath-TIR [23]
  • Coding: Persona Code [57], Evol CodeAlpaca [58]
  • Safety & Non-Compliance: CoCoNot [59], WildJailbreak [60], WildGuardMix [61]
  • Multilingual: Aya [62]
  • Precise Instruction Following: Persona IF [63]
@@ B.2 SmolTalk
The SmolTalk dataset was developed to address the lower instruction-tuned performance of the SmolLM2 base model [9]. Specifically, Allal et al. [9] blend high-quality conversational, task-specific, math, and code datasets, filtered and generated via Distilabel [64] annotations, to cover a wide range of instruction‐following capabilities. This results in a multi‐domain post-training corpus of 1,043,917 training samples which is used for SFT of SmolLM2 to boost instruction following, reasoning, and conversational skills in a reproducible, open-source pipeline. The SmolTalk subsets can be similarly grouped into seven high-level categories:
  • General: Everyday-Conversations [65], LongAlign [46], OpenHermes2.5 [12], Smol-Magpie-Ultra [9], Self-OSS-Starcoder-2-Instruct (Self-OSS-2) [66], SystemChats2.0 [67]
  • Knowledge Recall: Smol-Summarization [9]
  • Math: MetaMathQA-50k [24]
  • Reasoning: NuminaMath-CoT [23]
  • Coding: APIGen-80k [26]
  • Safety & Non-Compliance: Smol-Constraints [9]
  • Precise Instruction Following: Explore-Instruct-Rewriting [68], Smol-Rewrite [9]
@@ C.1.1 General Overview
Magpie [14] is a self-synthesis pipeline that extracts alignment annotations from open-weight, instruction-tuned LLMs without relying on seed prompts or human supervision. While Magpie can generate synthetic instruction-response pairs, we focus in this work on its annotation capabilities.
@@
Magpie is fully modular such that the judge model can be substituted by any LLM in principle. By default, Magpie uses Llama-3-8B-Instruct [27] for most annotation tasks, FsfairX-LLaMA3-RM-v0.1 [14] for instruct reward scoring, and Llama-Guard 2 [29] for safety classification.
@@ C.1.2 Extensions to Magpie
Multi-Turn Instruct Reward: While adapting most annotation tags is straightforward, computing instruct rewards for MT conversations is more complex. Magpie uses FsfairX-LLaMA3-RM-v0.1 [14], a reward model that assigns a continuous reward score $r^{*}$ to each instruction-response pair. To contextualize this score, it also computes a baseline reward $r_{\text{base}}$ using a reference model (typically the main LLM judge) on the same instruction. The difference $\Delta r=r^{*}-r_{\text{base}}$ reflects the relative improvement in response quality and is reported as the instruct reward.
@@ C.1.3 Choice of Judge Model
Magpie supports the use of any LLM as a potential annotation judge. In our experiments, we use Llama-3.3-70B-Instruct [27] as the primary judge model, based on two key considerations.
First, preliminary experiments with Qwen-based annotators revealed systematic biases. Fig. 8(a) and Fig. 8(b) show input quality distributions on 30k stratified subsets of Tulu and SmolTalk when annotated with Llama-3.3-70B-Instruct [27] versus Qwen2-72B-Instruct [69]. In both cases, Qwen strongly over-predicts the excellent label, while Llama produces more balanced annotations. In fact, our later analysis reveals a broader spread of input quality, with samples labeled as good, excellent, and even some rated as average or poor. Qwen tends to ignore these lower bands, particularly for SmolTalk, resulting in a less nuanced and potentially biased annotation profile.
@@ C.2.1 Token Length Distribution
To examine the token length distribution across the post-training datasets, we binned the per-sample token counts for both Tulu and SmolTalk (fine-tuned with Llama models) into 40 logarithmically spaced intervals ranging from $2^{4}$ (16) to $2^{13}$ (8,192) tokens. Fig. 9 and Fig. 10 show the resulting token length distributions across source subsets for Tulu and SmolTalk, respectively. Notably, different prompt sources exhibit distinct token length profiles, which can influence batch size, memory requirements, and learning dynamics when mixed during training. This variation in token lengths motivates our use of a sum-reduction over token-level losses, rather than the more commonly used mean-reduction (see App. E). A more in-depth discussion of this analysis is provided in Lambert et al. [7]. In our subsequent analysis, we do not further investigate token length characteristics, but include this section here for completeness.
@@ C.4.1 Overall Input Quality Distribution
Fig. 14(a) shows the overall distribution of input quality labels across all Tulu samples, while Fig. 14(b) breaks down the distribution by ST and MT samples. In addition, Table 10 reports the supporting statistics. Tulu is predominantly composed of high-quality inputs, with over 80% of samples rated as either “excellent” or “good”. This aligns with the strict curation and quality filtering practices described in Lambert et al. [7]. A similar trend holds for ST samples, which make up 95% of the dataset. For MT samples, the distribution is more balanced, with most samples rated as “good”, followed by “excellent”. However, a non-negligible portion of MT samples (26.5%) fall into the “poor” or “very poor” categories. As discussed in the main paper, this motivates a rigorous quality filtering step when constructing new data mixtures. This is further supported by our later analysis of instruct reward scores, which reveals a clear correlation between poor input quality and poor response quality.
@@
Similarly, Fig. 15(a) shows the overall distribution of input quality labels across all SmolTalk samples and Fig. 15(b) breaks down the distribution by ST and MT samples, with Table 11 reporting the statistics. SmolTalk also contains predominantly high-quality inputs, with 85% of samples rated as either “excellent” or “good”. In contrast to Tulu, significantly fewer samples fall into the “poor” or “very poor” categories, suggesting that Allal et al. [9] applied stricter quality control measures during curation. This trend holds across both ST and MT samples. In particular, the ST subset is even more skewed toward high-quality inputs, with over 95% of ST samples rated as “excellent” or “good”. These findings indicate that the SmolTalk data curation process results in consistently high-quality samples, across both turn types.
@@ C.5 Response Quality Analysis
Magpie rates the response quality, i.e., the quality of the assistant’s response to a user prompt, as the instruct reward via a language model-based reward model. Specifically, it employs FsfairX-LLaMA3-RM-v0.1 [70, 71, 72], a Llama-3-8B-Instruct-based reward model that assigns a continuous-valued score to each response. Since the original implementation supports only single-turn conversations, we extend the reward annotation pipeline to handle multi-turn samples by using a separate Llama-3.3-70B-Instruct-based LLM-as-a-Judge to evaluate MT responses on a scale between 0-5. We refer to more details in App. C.
@@ C.8.1 Overall Safety Distribution
Fig. 39 shows the overall safety distribution for the Tulu and SmolTalk datasets. In general, both datasets are predominantly safe, with 97.4% of Tulu samples and 99.1% of SmolTalk samples flagged as safe. This reflects the rigorous quality control applied by Lambert et al. [7] and Allal et al. [9], which includes an explicit focus on safety during data collection and filtering. Nevertheless, a small fraction of samples are labeled as unsafe. Below, we provide two representative examples from Tulu and SmolTalk.
@@
In this example, the assistant provides specialized advice, which is flagged by Llama-Guard 2 [29] as inappropriate, particularly due to its medical context. Even if the content appears benign or well-intentioned, Llama-Guard still classifies it under one of its 11 unsafe categories.
@@
These examples illustrate that most unsafe samples are benign in appearance but contain elements, such as medical or gambling-related content that trigger conservative safety filters. Nonetheless, the proportion of such flagged instances is negligible in both datasets. However, safety in LLM fine-tuning remains an active area of research [73], with various mechanisms available to instill safety both during and after training [74, 75, 76]. We leave a more comprehensive safety analysis of post-training datasets to future work.
@@ E.1 Fine-Tuning Configurations
To ensure reproducibility and comparability, we fine-tune all models using AllenAI’s Open-Instruct framework [^9] https://github.com/allenai/open-instruct, covering both SFT and DPO. By default, Open-Instruct applies a sum-reduction over token-level losses, rather than the more commonly used mean-reduction. This design choice ensures length-equitable weighting, where short and long sequences contribute proportionally to the total loss, preventing shorter examples from disproportionately influencing the gradient due to having fewer tokens. Moreover, summing losses leads to more stable optimization by avoiding fluctuations in loss scale caused by variation in batch composition or sequence length distributions. We refer to a more detailed analysis and discussion in [7].
@@ E.1.1 Supervised Fine-Tuning (SFT)
We fine-tune Llama-3.1-8B [27] and SmolLM2-1.7B [9] models on the Tulu [7], SmolTalk [9], Orca [16], and our proposed TuluTalk dataset. These two models are selected for consistency with prior work, being the default backbones in the respective dataset papers for Tulu and SmolTalk.
Fine-tuning is performed using BF16 mixed precision with Fully Sharded Data Parallelism (FSDP) on 8 × NVIDIA A100 80GB GPUs. To isolate dataset effects, we fix SFT hyperparameters per model across all experiments. For Llama-3.1-8B, we adopt the same hyperparameters as in Lambert et al. [7], and for SmolLM2-1.7B, we follow Allal et al. [13]. Table 14 provides the SFT training configurations for both models.
@@ E.1.2 Direct Preference Optimization (DPO)
For DPO, we choose the preference mixture [^10] https://huggingface.co/datasets/allenai/llama-3.1-tulu-3-8b-preference-mixture created by Lambert et al. [7], which is a curated blend of on-policy and off-policy preference data, synthetic instruction following augmentations, WildChat [8] conversational pairs, cleaned UltraFeedback [47] data, and a small Persona IF [63] subset, designed to balance broad performance and targeted instruction following under the DPO objective. We use the same hyperparameters from [7]. We set the KL-penalty coefficient (referred to as dpo_beta in Open-Instruct) to 5 and use the length-normalized DPO loss (dpo_loss_type=norm), following the recommendation in Lambert et al. [7]. We apply DPO on models that have already been fine-tuned via SFT using the Tulu, SmolTalk, and TuluTalk datasets. The full set of DPO hyperparameters is provided in Table 14.
@@ E.2 Evaluation Setup
We assess model performance using the LM Evaluation Harness framework [28], a widely adopted standard for evaluating language models across diverse benchmark suites. To ensure a comprehensive and task-diverse evaluation, we include benchmarks spanning Knowledge (e.g., MMLU [77], TruthfulQA [78]), Reasoning (e.g., BBH [79], ARC-C [80]), Commonsense Understanding (e.g., HellaSwag [81], WinoGrande [82]), Instruction Following (e.g., IF-Eval [83]), Mathematical Reasoning (e.g., GSM8K [84], MATH [85]), and Coding (e.g., HumanEval, HumanEval+ [86]). We further include benchmarks from Open LLM Leaderboards [17, 18] to gauge general instruction performance under competitive public standards. This setup ensures a fair, fine-grained comparison between models and data mixtures, highlighting both strengths and failure modes across capabilities.
@@ E.3.4 Performance Results for Diverse Models and Scales
To demonstrate the effectiveness and generalizability of TuluTalk across different architectures and scales, we provide results for three additional models: Qwen2.5-0.5B and Qwen2.5-3B [22], as well as for SmolLM3-3B [31], covering small- to mid-scale models of different architectures.
@@
To assess whether our SFT curation insights transfer to preference-tuned models, we also apply DPO on Llama models fine-tuned on Tulu, SmolTalk, and our proposed TuluTalk mixture. Table 19 reports DPO and SFT performance results across benchmarks for the Llama-3.1-8B base model, fine-tuned on all datasets under consideration and DPO-tuned using the same preference mixture proposed by Lambert et al. [7].

### v3

@@ B.1 Tulu
The Tulu dataset was created to bridge proprietary and open-source post-training data by leveraging publicly available datasets, persona-driven synthetic prompts, and rigorous decontamination procedures to mitigate test set leakage. Specifically, lambert2025tulu3pushingfrontiers collected 23,327,961 candidate prompts from over 20 distinct sources, curating a multi-skill SFT corpus that comprises 939,344 samples, forming the original Tulu-3-SFT-Mix data mixture. The Tulu subsets and their respective samples can be broadly categorized into nine high-level groups:
  • General: OpenAssistant (OASST1) [kopf2023openassistant], No Robots [no_robots], WildChat [zhao2024wildchat], UltraFeedback (Tülu HC-10) [cui2024ultrafeedbackboostinglanguagemodels]
  • Knowledge Recall: FLAN v2 [longpre2023flan], SciRIFF [wadden2024sciriffresourceenhancelanguage], TableGPT [tablegpt]
  • Math: Persona MATH, Persona MATH (Grade) [luo2025personamathboostingmathematicalreasoning]
  • Reasoning: Persona Algebra [persona_algebra], OpenMathInstruct2 [toshniwal2024openmathinstruct2acceleratingaimath], NuminaMath-TIR [numina_math_datasets]
  • Coding: Persona Code [persona_code], Evol CodeAlpaca [luo2024wizardcoder]
  • Safety & Non-Compliance: CoCoNot [deng2024coconutmodernizingcocosegmentation], WildJailbreak [wildteaming2024], WildGuardMix [wildguard2024]
  • Multilingual: Aya [singh2024ayadatasetopenaccesscollection]
  • Precise Instruction Following: Persona IF [persona_if]
@@ B.2 SmolTalk
The SmolTalk dataset was developed to address the lower instruction-tuned performance of the SmolLM2 base model [allal2025smollm2smolgoesbig]. Specifically, allal2025smollm2smolgoesbig blend high-quality conversational, task-specific, math, and code datasets, filtered and generated via Distilabel [distilabel-argilla-2024] annotations, to cover a wide range of instruction‐following capabilities. This results in a multi‐domain post-training corpus of 1,043,917 training samples which is used for SFT of SmolLM2 to boost instruction following, reasoning, and conversational skills in a reproducible, open-source pipeline. The SmolTalk subsets can be similarly grouped into seven high-level categories:
  • General: Everyday-Conversations [everydayconversations2024], LongAlign [bai2024longalign], OpenHermes2.5 [OpenHermes2.5], Smol-Magpie-Ultra [allal2025smollm2smolgoesbig], Self-OSS-Starcoder-2-Instruct (Self-OSS-2) [self_oss_2], SystemChats2.0 [syschats]
  • Knowledge Recall: Smol-Summarization [allal2025smollm2smolgoesbig]
  • Math: MetaMathQA-50k [yu2023metamath]
  • Reasoning: NuminaMath-CoT [numina_math_datasets]
  • Coding: APIGen-80k [liu2024apigen]
  • Safety & Non-Compliance: Smol-Constraints [allal2025smollm2smolgoesbig]
  • Precise Instruction Following: Explore-Instruct-Rewriting [wan2023explore], Smol-Rewrite [allal2025smollm2smolgoesbig]
@@ C.1.1 General Overview
Magpie [xu2024magpie] is a self-synthesis pipeline that extracts alignment annotations from open-weight, instruction-tuned LLMs without relying on seed prompts or human supervision. While Magpie can generate synthetic instruction-response pairs, we focus in this work on its annotation capabilities.
@@
Magpie is fully modular such that the judge model can be substituted by any LLM in principle. By default, Magpie uses Llama-3-8B-Instruct [grattafiori2024llama3herdmodels] for most annotation tasks, FsfairX-LLaMA3-RM-v0.1 [xu2024magpie] for instruct reward scoring, and Llama-Guard 2 [metallamaguard2] for safety classification.
@@ C.1.2 Extensions to Magpie
Multi-Turn Instruct Reward: While adapting most annotation tags is straightforward, computing instruct rewards for MT conversations is more complex. Magpie uses FsfairX-LLaMA3-RM-v0.1 [xu2024magpie], a reward model that assigns a continuous reward score $r^{*}$ to each instruction-response pair. To contextualize this score, it also computes a baseline reward $r_{\text{base}}$ using a reference model (typically the main LLM judge) on the same instruction. The difference $\Delta r=r^{*}-r_{\text{base}}$ reflects the relative improvement in response quality and is reported as the instruct reward.
@@ C.1.3 Choice of Judge Model
Magpie supports the use of any LLM as a potential annotation judge. In our experiments, we use Llama-3.3-70B-Instruct [grattafiori2024llama3herdmodels] as the primary judge model, based on two key considerations.
First, preliminary experiments with Qwen-based annotators revealed systematic biases. Fig. 8(a) and Fig. 8(b) show input quality distributions on 30k stratified subsets of Tulu and SmolTalk when annotated with Llama-3.3-70B-Instruct [grattafiori2024llama3herdmodels] versus Qwen2-72B-Instruct [yang2024qwen2technicalreport]. In both cases, Qwen strongly over-predicts the excellent label, while Llama produces more balanced annotations. In fact, our later analysis reveals a broader spread of input quality, with samples labeled as good, excellent, and even some rated as average or poor. Qwen tends to ignore these lower bands, particularly for SmolTalk, resulting in a less nuanced and potentially biased annotation profile.
@@ C.2.1 Token Length Distribution
To examine the token length distribution across the post-training datasets, we binned the per-sample token counts for both Tulu and SmolTalk (fine-tuned with Llama models) into 40 logarithmically spaced intervals ranging from $2^{4}$ (16) to $2^{13}$ (8,192) tokens. Fig. 9 and Fig. 10 show the resulting token length distributions across source subsets for Tulu and SmolTalk, respectively. Notably, different prompt sources exhibit distinct token length profiles, which can influence batch size, memory requirements, and learning dynamics when mixed during training. This variation in token lengths motivates our use of a sum-reduction over token-level losses, rather than the more commonly used mean-reduction (see App. E). A more in-depth discussion of this analysis is provided in lambert2025tulu3pushingfrontiers. In our subsequent analysis, we do not further investigate token length characteristics, but include this section here for completeness.
@@ C.4.1 Overall Input Quality Distribution
Fig. 14(a) shows the overall distribution of input quality labels across all Tulu samples, while Fig. 14(b) breaks down the distribution by ST and MT samples. In addition, Table 10 reports the supporting statistics. Tulu is predominantly composed of high-quality inputs, with over 80% of samples rated as either “excellent” or “good”. This aligns with the strict curation and quality filtering practices described in lambert2025tulu3pushingfrontiers. A similar trend holds for ST samples, which make up 95% of the dataset. For MT samples, the distribution is more balanced, with most samples rated as “good”, followed by “excellent”. However, a non-negligible portion of MT samples (26.5%) fall into the “poor” or “very poor” categories. As discussed in the main paper, this motivates a rigorous quality filtering step when constructing new data mixtures. This is further supported by our later analysis of instruct reward scores, which reveals a clear correlation between poor input quality and poor response quality.
@@
Similarly, Fig. 15(a) shows the overall distribution of input quality labels across all SmolTalk samples and Fig. 15(b) breaks down the distribution by ST and MT samples, with Table 11 reporting the statistics. SmolTalk also contains predominantly high-quality inputs, with 85% of samples rated as either “excellent” or “good”. In contrast to Tulu, significantly fewer samples fall into the “poor” or “very poor” categories, suggesting that allal2025smollm2smolgoesbig applied stricter quality control measures during curation. This trend holds across both ST and MT samples. In particular, the ST subset is even more skewed toward high-quality inputs, with over 95% of ST samples rated as “excellent” or “good”. These findings indicate that the SmolTalk data curation process results in consistently high-quality samples, across both turn types.
@@ C.5 Response Quality Analysis
Magpie rates the response quality, i.e., the quality of the assistant’s response to a user prompt, as the instruct reward via a language model-based reward model. Specifically, it employs FsfairX-LLaMA3-RM-v0.1 [fsfairX, xiong2024iterative, dong2023raft], a Llama-3-8B-Instruct-based reward model that assigns a continuous-valued score to each response. Since the original implementation supports only single-turn conversations, we extend the reward annotation pipeline to handle multi-turn samples by using a separate Llama-3.3-70B-Instruct-based LLM-as-a-Judge to evaluate MT responses on a scale between 0-5. We refer to more details in App. C.
@@ C.8.1 Overall Safety Distribution
Fig. 39 shows the overall safety distribution for the Tulu and SmolTalk datasets. In general, both datasets are predominantly safe, with 97.4% of Tulu samples and 99.1% of SmolTalk samples flagged as safe. This reflects the rigorous quality control applied by lambert2025tulu3pushingfrontiers and allal2025smollm2smolgoesbig, which includes an explicit focus on safety during data collection and filtering. Nevertheless, a small fraction of samples are labeled as unsafe. Below, we provide two representative examples from Tulu and SmolTalk.
@@
In this example, the assistant provides specialized advice, which is flagged by Llama-Guard 2 [metallamaguard2] as inappropriate, particularly due to its medical context. Even if the content appears benign or well-intentioned, Llama-Guard still classifies it under one of its 11 unsafe categories.
@@
These examples illustrate that most unsafe samples are benign in appearance but contain elements, such as medical or gambling-related content that trigger conservative safety filters. Nonetheless, the proportion of such flagged instances is negligible in both datasets. However, safety in LLM fine-tuning remains an active area of research [yao2024survey], with various mechanisms available to instill safety both during and after training [bianchi2024safetytunedllamaslessonsimproving, hsu2025safelorasilverlining, djuhera2025safemergepreservingsafetyalignment]. We leave a more comprehensive safety analysis of post-training datasets to future work.
@@ E.1 Fine-Tuning Configurations
To ensure reproducibility and comparability, we fine-tune all models using AllenAI’s Open-Instruct framework [^9] https://github.com/allenai/open-instruct, covering both SFT and DPO. By default, Open-Instruct applies a sum-reduction over token-level losses, rather than the more commonly used mean-reduction. This design choice ensures length-equitable weighting, where short and long sequences contribute proportionally to the total loss, preventing shorter examples from disproportionately influencing the gradient due to having fewer tokens. Moreover, summing losses leads to more stable optimization by avoiding fluctuations in loss scale caused by variation in batch composition or sequence length distributions. We refer to a more detailed analysis and discussion in [lambert2025tulu3pushingfrontiers].
@@ E.1.1 Supervised Fine-Tuning (SFT)
We fine-tune Llama-3.1-8B [grattafiori2024llama3herdmodels] and SmolLM2-1.7B [allal2025smollm2smolgoesbig] models on the Tulu [lambert2025tulu3pushingfrontiers], SmolTalk [allal2025smollm2smolgoesbig], Orca [mitra2024agentinstruct], and our proposed TuluTalk dataset. These two models are selected for consistency with prior work, being the default backbones in the respective dataset papers for Tulu and SmolTalk.
Fine-tuning is performed using BF16 mixed precision with Fully Sharded Data Parallelism (FSDP) on 8 × NVIDIA A100 80GB GPUs. To isolate dataset effects, we fix SFT hyperparameters per model across all experiments. For Llama-3.1-8B, we adopt the same hyperparameters as in lambert2025tulu3pushingfrontiers, and for SmolLM2-1.7B, we follow allal2024SmolLM. Table 14 provides the SFT training configurations for both models.
@@ E.1.2 Direct Preference Optimization (DPO)
For DPO, we choose the preference mixture [^10] https://huggingface.co/datasets/allenai/llama-3.1-tulu-3-8b-preference-mixture created by lambert2025tulu3pushingfrontiers, which is a curated blend of on-policy and off-policy preference data, synthetic instruction following augmentations, WildChat [zhao2024wildchat] conversational pairs, cleaned UltraFeedback [cui2024ultrafeedbackboostinglanguagemodels] data, and a small Persona IF [persona_if] subset, designed to balance broad performance and targeted instruction following under the DPO objective. We use the same hyperparameters from [lambert2025tulu3pushingfrontiers]. We set the KL-penalty coefficient (referred to as dpo_beta in Open-Instruct) to 5 and use the length-normalized DPO loss (dpo_loss_type=norm), following the recommendation in lambert2025tulu3pushingfrontiers. We apply DPO on models that have already been fine-tuned via SFT using the Tulu, SmolTalk, and TuluTalk datasets. The full set of DPO hyperparameters is provided in Table 14.
@@ E.2 Evaluation Setup
We assess model performance using the LM Evaluation Harness framework [lm-eval-harness], a widely adopted standard for evaluating language models across diverse benchmark suites. To ensure a comprehensive and task-diverse evaluation, we include benchmarks spanning Knowledge (e.g., MMLU [hendryckstest2021], TruthfulQA [lin-etal-2022-truthfulqa]), Reasoning (e.g., BBH [suzgun2022challenging], ARC-C [Clark2018ThinkYH]), Commonsense Understanding (e.g., HellaSwag [zellers2019hellaswag], WinoGrande [sakaguchi2019winogrande]), Instruction Following (e.g., IF-Eval [zhou2023instructionfollowing]), Mathematical Reasoning (e.g., GSM8K [cobbe2021training], MATH [hendrycksmath2021]), and Coding (e.g., HumanEval, HumanEval+ [chen2021codex]). We further include benchmarks from Open LLM Leaderboards [leaderboardv1, open-llm-leaderboard-v2] to gauge general instruction performance under competitive public standards. This setup ensures a fair, fine-grained comparison between models and data mixtures, highlighting both strengths and failure modes across capabilities.
@@ E.3.4 Performance Results for Diverse Models and Scales
To demonstrate the effectiveness and generalizability of TuluTalk across different architectures and scales, we provide results for three additional models: Qwen2.5-0.5B and Qwen2.5-3B [yang2024qwen2], as well as for SmolLM3-3B [bakouch2025smollm3], covering small- to mid-scale models of different architectures.
@@
To assess whether our SFT curation insights transfer to preference-tuned models, we also apply DPO on Llama models fine-tuned on Tulu, SmolTalk, and our proposed TuluTalk mixture. Table 19 reports DPO and SFT performance results across benchmarks for the Llama-3.1-8B base model, fine-tuned on all datasets under consideration and DPO-tuned using the same preference mixture proposed by lambert2025tulu3pushingfrontiers.